OpenStudy (anonymous):

let x be the number of heads in two tosses of a biased coin with P(H)=.4. Work out the mean and variance of X?

OpenStudy (anonymous):

It's a binomial distribution, right?

OpenStudy (anonymous):

That would make the expected value \(np\) and the variance \(np(1-p)\)

OpenStudy (anonymous):

In this case \(n=2\) and \(p=0.4\). Do you want to do this manually?

OpenStudy (anonymous):

yes, please

OpenStudy (anonymous):

Okay, so there are only \(2^2=4\) outcomes.\[ HH, HT, TH, TT \]

OpenStudy (anonymous):

Remember that \(P(T)= 1-P(H) = 1-0.4=0.6\). The respective probabilities are: \[ (0.4)(0.4), (0.4)(0.6), (0.6)(0.4),(0.6)(0.6) \]

OpenStudy (anonymous):

Do you understand these probabilities? Remember the flips are independent events so probability of getting heads twice is just \(\Pr(H)\times \Pr(H)\)

OpenStudy (anonymous):

X=2 (prob=1/4) x=1 (prob=1/2) x=0 (prob=1/4)

OpenStudy (anonymous):

oh. wait, i got how you got those probabilities. but how did you get 0.6?

OpenStudy (anonymous):

\(1-0.4=0.6\)

OpenStudy (anonymous):

got it! how do i solve it manually?

OpenStudy (anonymous):

The respective value of \(x\) for these outcomes are \[ 2, 1, 1, 0 \]

OpenStudy (anonymous):

The expected value is: The sum of: The probability of each outcome times the value of that outcome

OpenStudy (anonymous):

So \[ E[X] = 2(0.4)(0.4)+1(0.4)(0.6)+1(0.6)(0.4)+0(0.6)(0.6) \]

OpenStudy (anonymous):

Which simplifies to \(0.8\). And checking it with the typical binomial distribution: \(n=2,p=0.4\): \[2\times 0.4=0.8\]

OpenStudy (anonymous):

ok, that's what i got the first time! thank you!

OpenStudy (anonymous):

For the variance, we have to use different values. We subtract the expected value from the normal value and square it.

OpenStudy (anonymous):

so i would use the .16, .24, .24, .16 ?? multiplying it with 2, 1, 1, and 0?

OpenStudy (anonymous):

I'm not sure where you are getting those numbers.

OpenStudy (anonymous):

the probabilities of 0.4*0.4=.16???

OpenStudy (anonymous):

The values \[ 2,1,1,0 \]Need to be replaced with \[ (2-0.8)^2,(1-0.8)^2,(1-0.8)^2,(0-0.8)^2 \]

OpenStudy (anonymous):

\((0.4)(0.4)=0.16\) But \((0.6)(0.6)=0.36\)

OpenStudy (anonymous):

right! (2*.16) + (1*.24) + (1*.24) + (0*.36)

OpenStudy (anonymous):

Yes, that is for the expected value.

OpenStudy (anonymous):

that would the sigma^2 answer correct?

OpenStudy (anonymous):

how do i find the variance?

OpenStudy (anonymous):

It's almost the same as expected value except for each value, you change it to \((x-\mu)^2\) The values \[ 2,1,1,0 \]Need to be replaced with \[ (2-0.8)^2,(1-0.8)^2,(1-0.8)^2,(0-0.8)^2 \]

OpenStudy (anonymous):

why are you subtracting it by .8?

OpenStudy (anonymous):

i can't get it to equal .48...

OpenStudy (anonymous):

Because \(0.8\) is the expected value.

OpenStudy (anonymous):

\[ (2−0.8)^2(0.4)(0.4)+(1−0.8)^2(0.4)(0.6)+(1−0.8)^2(0.6)(0.4)+(0−0.8)^2(0.6)(0.6) \]

OpenStudy (anonymous):

That comes out to \(0.48\) for me.